OpenStudy (anonymous):

You have 16 coins that are heads up and 18 coins that are tails up. After you add some coins that are tails up, the ratio of heads up to tails up is 1 to 1.5 How many coins are tails up? How many coins in all?

jimthompson5910 (jim_thompson5910):

Let x = # of coins you're adding that are tails up the initial fraction of heads to tails is 16/18 if you add in x more coins that are tails up, then you get 16/(18+x) This fraction is equal to 1/1.5, which means 16/(18+x) = 1/1.5

OpenStudy (anonymous):

Whoops, read the question wrong... Amendment: h = heads-up coins t = tails-up coins h/t = 16/18 at first then: h/t = 1/1.5 = 16/x => 2/3 = 16/x Solve for x.

OpenStudy (anonymous):

How many total coins

jimthompson5910 (jim_thompson5910):

There will be 16+18+x total coins

jimthompson5910 (jim_thompson5910):

after you solve 16/(18+x) = 1/1.5 for x, you use it to evaluate 16+18+x to get the total